Iron metal (Fe) reacts with chlorine gas (Cl2). How many grams produced win 515 g of Cl2 reacts with excess Fe. considering molar mass CLT is 70.90 g/ mole , and the molar mass of FeCL3 is 162.2 g/mole
Answer #1

To determine the number of grams of FeCl3 produced when 515 g of Cl2 reacts with excess Fe, we need to use the balanced chemical equation for the reaction between Fe and Cl2:

2 Fe + 3 Cl2 -> 2 FeCl3

From the equation, we can see that the ratio between Cl2 and FeCl3 is 3:2. This means that for every 3 moles of Cl2, we produce 2 moles of FeCl3.

First, we need to calculate the number of moles of Cl2 in 515 g:

Molar mass of Cl2 = 2 * molar mass of Cl = 2 * 35.45 g/mol = 70.90 g/mol

Number of moles of Cl2 = mass of Cl2 / molar mass of Cl2 = 515 g / 70.90 g/mol ≈ 7.26 moles

Since the ratio of Cl2 to FeCl3 is 3:2, the number of moles of FeCl3 produced will be:

Number of moles of FeCl3 = (2/3) * number of moles of Cl2 = (2/3) * 7.26 moles ≈ 4.84 moles

Finally, we can calculate the mass of FeCl3 produced using its molar mass:

Mass of FeCl3 = number of moles of FeCl3 * molar mass of FeCl3 = 4.84 moles * 162.2 g/mol ≈ 785.53 g

Therefore, approximately 785.53 grams of FeCl3 will be produced when 515 grams of Cl2 reacts with excess Fe.

answered by: Hydra Master